# The ambiguous AT2022rze: Changing-look AGN mimicking a supernova in a merging galaxy system

### v1

@@ Appendix A Light curves
58205.28 | 0.18 | 5.46 | 20.31 | ZTFg | T | |
58214.25 | -1.38 | 12.54 | 19.41 | ZTFg | T | |
58216.25 | -0.26 | 4.27 | 20.58 | ZTFg | T |
@@
60007.18 | 19.00 | 0.32 | 18.59 | 91.27 | 26.55 | sdssr |
60012.32 | 18.98 | 0.07 | 20.15 | 93.25 | 6.31 | sdssr |
60012.35 | 19.44 | 60.96 | sdssg | |||
60012.35 | 18.10 | 209.65 | sdssr | |||
60071.33 | 18.66 | 0.15 | 19.05 | 124.36 | 17.36 | sdssr |
60071.36 | 19.53 | 55.87 | sdssg | |||
60071.36 | 18.66 | 0.09 | 19.62 | 124.87 | 10.26 | sdssr |
60071.36 | 17.88 | 0.05 | 19.39 | 255.08 | 12.75 | sdssi |
60263.52 | 18.86 | 0.07 | 20.14 | 103.30 | 6.40 | sdssr |
60378.20 | 20.02 | 0.10 | 20.82 | 35.72 | 3.42 | sdssg |
60378.21 | 19.77 | 0.06 | 21.08 | 45.06 | 2.68 | sdssr |
60378.21 | 18.69 | 0.04 | 20.65 | 120.96 | 3.98 | sdssi |
60525.17 | 19.79 | 0.09 | 20.79 | 44.25 | 3.50 | sdssr |
60525.17 | 18.83 | 0.05 | 20.41 | 106.88 | 4.97 | sdssi |
60529.17 | 19.80 | 0.07 | 21.05 | 43.53 | 2.75 | sdssr |
60529.17 | 18.78 | 0.04 | 20.75 | 111.94 | 3.63 | sdssi |
60535.16 | 20.69 | 19.23 | sdssr | |||
60535.16 | 18.77 | 0.04 | 20.65 | 112.63 | 4.00 | sdssi |
60730.54 | 21.20 | 12.04 | sdssr | |||
60730.55 | 19.42 | 0.07 | 20.70 | 61.95 | 3.82 | sdssi |

### v2

@@ Appendix A Light curves
| 58205.28 | 0.18 | 5.46 | 20.31 | ZTFg | T | |
| 58214.25 | -1.38 | 12.54 | 19.41 | ZTFg | T | |
| 58216.25 | -0.26 | 4.27 | 20.58 | ZTFg | T |
@@
| 60007.18 | 19.00 | 0.32 | 18.59 | 91.27 | 26.55 | sdssr |
| 60012.32 | 18.98 | 0.07 | 20.15 | 93.25 | 6.31 | sdssr |
| 60012.35 | 19.44 | 60.96 | sdssg | |||
| 60012.35 | 18.10 | 209.65 | sdssr | |||
| 60071.33 | 18.66 | 0.15 | 19.05 | 124.36 | 17.36 | sdssr |
| 60071.36 | 19.53 | 55.87 | sdssg | |||
| 60071.36 | 18.66 | 0.09 | 19.62 | 124.87 | 10.26 | sdssr |
| 60071.36 | 17.88 | 0.05 | 19.39 | 255.08 | 12.75 | sdssi |
| 60263.52 | 18.86 | 0.07 | 20.14 | 103.30 | 6.40 | sdssr |
| 60378.20 | 20.02 | 0.10 | 20.82 | 35.72 | 3.42 | sdssg |
| 60378.21 | 19.77 | 0.06 | 21.08 | 45.06 | 2.68 | sdssr |
| 60378.21 | 18.69 | 0.04 | 20.65 | 120.96 | 3.98 | sdssi |
| 60525.17 | 19.79 | 0.09 | 20.79 | 44.25 | 3.50 | sdssr |
| 60525.17 | 18.83 | 0.05 | 20.41 | 106.88 | 4.97 | sdssi |
| 60529.17 | 19.80 | 0.07 | 21.05 | 43.53 | 2.75 | sdssr |
| 60529.17 | 18.78 | 0.04 | 20.75 | 111.94 | 3.63 | sdssi |
| 60535.16 | 20.69 | 19.23 | sdssr | |||
| 60535.16 | 18.77 | 0.04 | 20.65 | 112.63 | 4.00 | sdssi |
| 60730.54 | 21.20 | 12.04 | sdssr | |||
| 60730.55 | 19.42 | 0.07 | 20.70 | 61.95 | 3.82 | sdssi |